OpenStudy (geometry_hater):

Which is the center of the circle whose equation is (x - 3)2 + (y - 4)2 = 121?

OpenStudy (geometry_hater):

I dont know how to find centers of circles

OpenStudy (amistre64):

(x-Cx)^2 + (y-Cy)^2 = r^2

OpenStudy (amistre64):

its not so much how to find it, but how to retrieve them from the equation

OpenStudy (amistre64):

(Cx,Cy) is the center

OpenStudy (geometry_hater):

Ok i still dont know what to do for the problem

OpenStudy (amistre64):

me either ....

OpenStudy (agreene):

\[ (x - 3)^2 + (y - 4)^2 = 121\] if that is your equation then the center is at: (3,4) because it is in general form, like amistre64 stated: \[(x-Cx)^2+(y-Cy)^2=r^2\] where Cx and Cy are your central coordinates (note the negatives in the general form, meaning if it were (x+3)^2+(y+4)^2 it would be centered at (-3,-4))

OpenStudy (geometry_hater):

so -3 = the x's and -4 equals the y's???

OpenStudy (agreene):

No, it is asking for the center of the circle. If the equation is the same as the one I posted then the center is (3,4) because the equation is in general form, and you can just look at the equation and get that information.